Required information Exercise 5-16 Break-Even Analysis and CVP Graphing [LO5-2, LO5-4, LO5-5] The following information applies to the questions displayed below.] The Hartford Symphony Guild is planning its annual dinner-dance. The dinner-dance committee has assembled the following expected costs for the event: Dinner (per person) Favors and program (per person) Band Rental of ballroom Professional entertainment during intermission Tickets and advertising $ 23 $ 780 $ 390 260 $1,170 The committee members would like to charge $40 per person for the evenings activities Exercise 5-16 Part 1 Required 1. What is the break-even point for the dinner-dance (in terms of the number of persons who must attend)? 2. Assume that last year only 100 persons attended the dinner-dance. If the same number attend this year, what price per ticket must be charged in order to break even? 1. Break-even point persons 2. Ticket price per person to break even
Answer #1

Requirement 1:

Fixed cost= band+ rental of ballroom+professional entertainment+ tickets and advertising

Note: the above expenses have been included in fixed costs because they do not change with the number of people attending,ie , they do not change with change in activity level for a given the range.

Fixed cost= 780+390+260+1170=$2600

Variable cost per unit= dinner per person+ favours and program per person

= 23+ 4= $27 per person

Ticket charge= $40 per person

Contribution per unit= ticket charge-variable cost per unit

=40-27=$13 per person

Break even point (units)=fixed cost/contribution per unit

Break even point (persons)=2600/13=200 persons

Requirement 2:

assuming break even point is 100 persons

break even point (persons)= fixed cost/contribution per unit

100= 2600/contribution per unit

Contribution per unit= 2600/100= $26 per unit

Ticket price per person to break even= new contribution per unit+ variable cost per unit= 26+27= $53 per person.